5x=3
x=-5/3
is this pair of lines parallel, perpendicular, or neither

Answer by Boreal(15235)

x=(3/5), vertical line at x=(3/5)
x=-5/3, vertical line at x=-5/3
These vertical lines are parallel to each other.